OpenStudy (anonymous):

solve the logarithmic equation rounding to the nearest ten-thousandth 3 log 2x = 4

OpenStudy (anonymous):

Divide both sides by 3 to start off

OpenStudy (anonymous):

Which is it, a 4 or a 1?

OpenStudy (anonymous):

Your original problem says 4 on the right side, but when you divided by 3 you said 1/3

OpenStudy (anonymous):

you said divide both sides by 3

OpenStudy (anonymous):

Right, you should have gotten log 2x=4/3

OpenStudy (anonymous):

Now, remember what we do to cancel the log?

OpenStudy (anonymous):

not sure what i did wrong

OpenStudy (anonymous):

How do we cancel the log?

OpenStudy (anonymous):

divide by 4/3

OpenStudy (anonymous):

No, what we did in the other problem.... remember \[10^{\log 2x}=10^{4/3}\]

OpenStudy (anonymous):

i keep forgetting that

OpenStudy (anonymous):

Burn it into ya brain. so now we have \[2x=10^{4/3}\] x=?

OpenStudy (anonymous):

Solve it like any other equation and once you have x alone plug it into a calculator \[x=10^{4/3}/2\]

OpenStudy (anonymous):

Plugging that into a calc should give you 10.772

OpenStudy (anonymous):

have a ti30x and not sure how to use it

OpenStudy (anonymous):

Make sure you use parentheses type it in like this (10^(4/3))/2

OpenStudy (anonymous):

Not sure whats wrong. Mine tells me 10.772

OpenStudy (anonymous):

i got it missed the ( in front of 4

OpenStudy (anonymous):

Not a problem